# Binarized Simplicial Convolutional Neural Networks

### v1

@@ 5.1 Citation Complex
The raw data of the Citation Complex is from Semantic Scholar Open Research Corpus [32], where the number of citations of $k+1$ authors forms a $k$-simplex where $k=0...5$. The 6 simplices naturally form a simplicial complex because they form a hierarchical structure: each $k$-simplex is a subset author collaborations of the $k+1$-simplex. To examine the performance of our Bi-SCNN algorithm, we will follow the training and testing procedure seen in SNN [15] and SCNN [21]. For each $k$-simplex a portion of the simplicial data will be missing at random. In this transductive learning task, the goal is to impute the missing data from the existing data. We will be testing 5 different missing rates: 10%, 20%, 30%, 40%, and 50%.
A 2-layer Bi-SCNN (Bi-SCNN-2) and a 3-layer Bi-SCNN (Bi-SCNN-3) will be compared against a 2-layer SNN (SNN-2), a 2-layer SCNN (SCNN-2), and a 3-layer SCNN (SCNN-3). The number of simplicial filters per layer is set to 30 for all algorithms. Each layer in SNN-2 and SCNN-2 will have a length $J=1$ to keep the model size roughly the same order of magnitude as the Bi-SCNN-2. SCNN-3 will have a filter length $J=2$, which is the best-performing setting seen in the original paper [21]. GCN is not included because SNN is equivalent to the GCN for $k=0$.
@@
Figure 5: Run time of each algorithm on the Citation Complex with $10\%$ missing data.
@@
Figure 6: Training loss of each algorithm on the Citation Complex with $10\%$ missing data.
@@
Looking at Figure 6, we can see that all the tested algorithms converge and their convergence behaviors are distinctively different. The different convergence behavior is an indication of them having different simplicial representations. From Table 1, we can see that 2-layer Bi-SCNN performs the best in most of the simplices under all the tested missing rates. This confirms that even though the Bi-SCNN uses a simpler propagation scheme by only having a length-1 simplicial filter, the embedding power is equivalent to the SNN or SCNN in the current experiment. Another interesting observation is that the 2-layer models perform slightly better than the 3-layer models in many cases. We think this is due to the fact that the 3-layer models have too many simplicial convolutional filters and the simplicial embedding got over-smoothed during the propagation process. If we look at the ratio of parameter size to run time, we can see that the Bi-SCNN is the most efficient among all the tested algorithms. Bi-SCNN-2 has 20% more parameters than SNN-2 but takes only $1/5$ the time that SNN needs to complete the experiment. We should point out that for roughly similar model sizes, the Bi-SCNN has higher accuracy in imputing the citations compared to the SNN because Bi-SCNN uses assigns weights separately by considering upper and lower simplicial adjacencies separately but SNN does not, making the Bi-SCNN more expressive. The Bi-SCNN is indeed a more concise model than the SNN and the SCNN because not only does Bi-SCNN use fewer parameters but also operates faster. The reason that the Bi-SCNN is time-efficient is that the binary-sign propagation takes fewer computations, which is proven in traditional CNN [33] and has also been proven when data is purely on the nodes of the graph by the G-Sign [28] and Bi-GCN [27]. We can conclude that the Bi-SCNN can effectively and efficiently impute the missing citations compared to the SCNN and SAT.
SAT SCNN Bi-SCNN
Id, 2 60.25$\pm$6.47 64.00$\pm$3.39 71.25$\pm$4.0
Id, 3 64.75$\pm$5.30 69.00$\pm$2.78 71.00 $\pm$4.06
LR, 2 61.75$\pm$7.24 77.00$\pm$4.00 79.50$\pm$3.67
LR, 3 62.25$\pm$2.08 81.75$\pm$2.97 77.00$\pm$4.15
Tanh, 2 64.00$\pm$6.04 74.75$\pm$6.75 78.75$\pm$5.94
Tanh, 3 66.5$\pm$5.50 75.25$\pm$6.66 73.75$\pm$3.01
@@
Table 3: Classification (test) accuracy on the Ocean Drifter Dataset. The number after the activation denotes how many simplicial layers.
SAT SCNN Bi-SCNN
Id, 2 1196 s 2197 s 143 s
Id, 3 1515 s 6450 s 232 s
LR, 2 1116 s 2187 s 165 s
LR, 3 1385 s 6441 s 231 s
Tanh, 2 1161 s 2196 s 150 s
Tanh, 3 1491 s 6399 s 234 s
Table 4: Experiment run time (in seconds) on the Ocean Drifter Dataset. The number after the activation denotes how many simplicial layers.
@@ 5.2 Ocean Drifter Complex
The Ocean Drifter Complex is a dataset formed by using simplicial embedding to represent real-world ocean drifters around Madagascar Island [14]. The ocean that surrounds Madagascar Island is divided into hexagonal grids and each grid is represented by a node in the 00-simplex. If the trajectory of an ocean drifter passes two adjacent hexagonal grids an edge will be formed between the node. In the dataset, there are two types of trajectories and the goal is to classify which type of trajectory the input trajectory belongs to. A visualization of this dataset is shown in Figure 4. We should emphasize that the formed simplex structure itself is undirected, but the data that resides on the simplex has orientations.
@@
The data provided 160 training trajectories and 40 testing trajectories all on the $1$-simplex. In other words, the feature only exists on the edges. Our baseline algorithms are SAT [22] and SCNN [21]. We will follow the batch training procedure seen in SAT [22] but with the batch size set to 40 and each experiment run consists of 5000 epochs. For all three algorithms, we will be running two different architectural settings: 2 simplicial layers and 3 simplicial layers. As suggested by [22], the simplicial layers are followed by a mean pooling layer, then 2 MLP readout layers, and finally a softmax layer. For each architectural setting, there will be three different choices of activation function: identity (id) function, leaky ReLu (LR) function, and tanh function. The SAT suggests the choices of Id and Tanh functions because odd activation functions lead to orientation-equivariant output, which is shown to perform well on a synthetic ocean drifter dataset [22]. The classification accuracy is shown in Table 3 and the time to complete the experiments is recorded in Table 4. The number of trainable parameters of each setting is shown in Table 5.
@@
In the Ocean Drifter Dataset, we can observe again that the Bi-SCNN has similar classification accuracy as the SCNN for all three choices of action functions. The reason Bi-SCNN is able to consistently outperform SAT is due to the fact that the Bi-SCNN utilizes the Hodge decomposition: Bi-SCNN optimizes parameters for upper and lower adjacencies separately but the SAT uses the standard Hodge Laplacian. Additionally, the processing of the harmonic component of the input simplicial features is absent in SAT, making its simplicial embedding less representative than convolution-based algorithms. One may notice that the run time reduction of Bi-SCNN is not as prominent in the ocean drifter complex as in the citation complex. This is because when we recorded the run time, the global pooling and the 2 MLP layers also contributed to the time in addition to the time it takes to run the simplicial layers. The Bi-SCNN does not optimize the global pooling and the 2 MLP layers, so these layers will perform indifferently among Bi-SCNN, SAT, and SCNN. Still, Bi-SCNN is the fastest among Bi-SCNN, SAT, and SCNN when all three are compared under the same number of simplicial layers.
Citation Ocean Drifter
SAT-2 - 5762
SAT-3 - 6722
SNN-2 906 -
SCNN-2 1986 4712
SCNN-3 29166 10142
Bi-SCNN-2 1146 3750
Bi-SCNN-3 15726 6450
Table 5: The number of trainable parameters of each network setting on both data sets.
@@ 6 Conclusion
We introduced Bi-SCNN, combining simplicial convolution with binary-sign propagation. Incorporating the upper and lower adjacencies using the Hodge Laplacian matrix, the Bi-SCNN can effectively capture higher-order structured simplicial features. Bi-SCNN outperforms previously proposed SNN algorithms on citation and ocean drifter complexes in terms of time efficiency, while maintaining comparable prediction accuracy.

### v2

@@ 5.1 Citation Complex
The raw data of the Citation Complex is from Semantic Scholar Open Research Corpus (Ammar et al.,, 2018), where the number of citations of $k+1$ authors forms a $k$-simplex where $k=0...5$. The 6 simplices naturally form a simplicial complex because they form a hierarchical structure: each $k$-simplex is a subset author collaborations of the $k+1$-simplex. The number of data elements in each simplicial order for $k=0...5$ is $N_{0}=352$, $N_{1}=1474$, $N_{2}=3285$, $N_{3}=5019$, $N_{4}=5559$, and $N_{5}=4547$. It is worth mentioning that each simplicial complex standing alone by itself could be viewed as a dataset as well. For example $\mathcal{S}_{0}$ is a graph with data on the nodes only, which means that GNNs such as the GAT and the GCN can be deployed on it. However, graph node algorithms are only suitable for data defined on the nodes, and for higher-order structures like the data on the graph edges or triangles formed by the edges, we still need to rely on simplicial complex algorithms.
To examine the performance of our Bi-SCNN algorithm, we will follow the training and testing procedure seen in SNN (Ebli et al.,, 2020) and SCNN (Yang et al., 2022a, ). For each $k$-simplex a portion of the simplicial data will be missing at random. In this transductive learning task, the goal is to impute the missing data from the existing data. We will be testing 5 different missing rates: 10%, 20%, 30%, 40%, and 50%.
A 2-layer Bi-SCNN (Bi-SCNN-2) and a 3-layer Bi-SCNN (Bi-SCNN-3) will be compared against a 2-layer SNN (SNN-2), a 2-layer SCNN (SCNN-2), a 3-layer SCNN (SCNN-3), and a 2-layer Simplicial MPNN (MPNN-2). The number of simplicial filters per layer is set to 30 for all algorithms except for MPNN-2. Each layer in SNN-2 and SCNN-2 will have a length $J=1$ to keep the model size roughly the same order of magnitude as the Bi-SCNN-2. SCNN-3 will have a filter length $J=2$, which is the best-performing setting seen in the original paper (Yang et al., 2022a, ). GCN is not included because SNN is equivalent to the GCN for $k=0$. Due to the memory demand of the MPNN-2 and the hardware constraint of our experiment machine, MPNN-2 will have 2 layers but each layer only has 4 simplicial filters to avoid out-of-memory (OOM) error. In theory, this will not affect the MPNN by much as we can represent the combination of $J$ filters using a single filter using TSP techniques: $\sum_{j=1}^{J}\mathbf{U}_{k}h^{(j)}(\mathbf{\Lambda}_{k})\mathbf{U}_{k}^{T}=\mathbf{U}_{k}\mathbf{\Sigma}_{k}\mathbf{U}_{k}^{T}$, where $\mathbf{\Sigma}_{k}=\sum_{j=1}^{J}h^{(j)}(\mathbf{\Lambda}_{k})$.
@@
Model Trainable Parameter
SNN-2 906
SCNN-2 1986
SCNN-3 29166
Bi-SCNN-2 1146
Bi-SCNN-3 15726
MPNN-2 503463946
Table 3: The number of trainable parameters for each model on the Citation dataset.
@@
Figure 8: The mean training loss for each algorithm on the Citation Complex under $10\%$ missing rate.
First, analyzing the accuracy in Table 1, we can see that SNN, SCNN, and Bi-SCNN have similar performances. This is because all three algorithms are based on simplicial convolutions. However, further analyzing Table 1, we can see that 2-layer Bi-SCNN performs the best in most of the simplices under all the tested missing rates. This confirms that even though the Bi-SCNN uses a simpler propagation scheme by only having a length-1 simplicial filter, the embedding power is equivalent to the SNN or SCNN in the current experiment. We think this is due to the fact that the 3-layer models have too many simplicial convolutional filters and the simplicial embedding got over-smoothed during the propagation process.
The performance of the MPNN is suboptimal due to the extensive number of trainable parameters, necessitating additional preprocessing techniques or implementation tricks to effectively train the architecture. Analyzing the output of MPNN, we found out that for $\mathcal{S}_{0}$ and $\mathcal{S}_{5}$, the MPNN over-smoothed the output: the majority of the output shares a similar value. This can be again confirmed by the low standard deviation of the accuracies. Another interesting observation is that the 2-layer models perform slightly better than the 3-layer models in many cases. As for $\mathcal{S}_{1}$, $\mathcal{S}_{2}$, $\mathcal{S}_{3}$, and $\mathcal{S}_{4}$, the high standard deviation can be caused by the following 3 reasons: 1. the parameter size and model complexity is too high compared to the data size; 2. the training duration is relatively too short compared to the parameter size; 3. poor initialization of parameters.
Looking at Figure 8, we can see that all the tested algorithms converge and their convergence behaviors are distinctively different. The different convergence behavior is an indication that SNN, SCNN, and Bi-SCNN have different simplicial representations. For MPNN-2, the relatively high loss can be attributed to an excessive number of trainable parameters, which makes MPNN difficult to be properly trained for only 4 filters, leading to high training loss.
@@
If we look at the ratio of parameter size to run time, we can see that the Bi-SCNN is the most efficient among all the tested algorithms. Bi-SCNN-2 has 20% more parameters than SNN-2 but takes only $1/5$ the time that SNN needs to complete the experiment. We should point out that for roughly similar model sizes, the Bi-SCNN has higher accuracy in imputing the citations compared to the SNN because Bi-SCNN assigns weights separately by considering upper and lower simplicial adjacencies separately but SNN does not, making the Bi-SCNN more expressive. The Bi-SCNN is indeed a more concise model than the SNN and the SCNN because not only does Bi-SCNN use fewer parameters but also operates faster. The reason that the Bi-SCNN is time-efficient is that the binary-sign propagation takes fewer computations, which is proven in traditional CNN (Courbariaux et al.,, 2016) and has also been proven when data is purely on the nodes of the graph by the G-Sign (Yan et al.,, 2022) and Bi-GCN (Wang et al., 2021b, ). We can conclude that the Bi-SCNN can effectively and efficiently impute the missing citations compared to the MPNN, SNN, and SCNN.
Id LR Tanh
SAT-2 60.25$\pm$6.47 61.75$\pm$7.24 64.00$\pm$6.04
SAT-3 64.75$\pm$5.30 62.25$\pm$2.08 66.50$\pm$5.50
SCNN-2 64.00$\pm$3.39 77.00$\pm$4.00 74.75$\pm$6.75
SCNN-3 69.00$\pm$2.78 81.75$\pm$2.97 75.25$\pm$6.66
Bi-SCNN-2 71.25$\pm$4.00 79.50$\pm$3.67 78.75$\pm$5.94
Bi-SCNN-3 71.00$\pm$4.06 77.00$\pm$4.15 73.75$\pm$3.01
MPNN* 46.5$\pm$5.7 73.0$\pm$2.7 72.5$\pm$0.0
@@
Table 4: Classification (test) accuracy on the Ocean Drifter Dataset. The number after the activation denotes how many simplicial layers. (Bold denotes best-performing; underline denotes second-to-the-best. MPNN results are obtained from (Bodnar et al.,, 2021)).
Id LR Tanh
SAT-2 1196.90 $\pm$ 114.80 s 1116.93 $\pm$ 51.48 s 1161.90 $\pm$ 28.95 s
SAT-3 1515.73 $\pm$ 5.97 s 1385.93 $\pm$ 1.77 s 1491.28 $\pm$ 11.58 s
SCNN-2 2197.16 $\pm$ 8.93 s 2187.98 $\pm$ 1.12 s 2196.65 $\pm$ 2.07 s
SCNN-3 6451.00 $\pm$ 6.47 s 6441.42 $\pm$ 10.62 s 6399.48 $\pm$ 2.83 s
Bi-SCNN-2 149.33 $\pm$ 0.87 s 151.06 $\pm$ 0.28 s 150.21 $\pm$ 0.50 s
Bi-SCNN-3 231.91 $\pm$ 0.84 s 232.54 $\pm$ 0.46 s 234.99 $\pm$ 0.19 s
MPNN-2 (ours) 1646.23 $\pm$ 0.30 s 1653.04 $\pm$ 2.08 s 1644.69 $\pm$ 3.12 s
Table 5: Experiment run time (in seconds) on the Ocean Drifter Dataset.
@@ 5.2 Ocean Drifter Complex
The Ocean Drifter Complex dataset utilizes simplicial embedding to represent real-world ocean drifters around Madagascar Island (Schaub et al.,, 2020). The ocean surrounding Madagascar Island is segmented into hexagonal grids, with each grid represented as a node in the 00-simplex. If the trajectory of an ocean drifter passes two adjacent hexagonal grids an edge will be formed between the node. In the dataset, there are two types of trajectories and the goal is to classify which type of trajectory the input trajectory belongs to. The simplicial complex of the dataset contains $N_{0}=133$ nodes, $N_{1}=320$ edges, and $N_{2}=186$ triangles; the simplicial complex is shared among all the trajectories. A visualization of this dataset is shown in Figure 9. We should emphasize that the formed simplex structure itself is undirected, but the data that resides on the simplex has orientations. The data provided 160 training trajectories and 40 testing trajectories all on the $1$-simplex. In other words, the feature only exists on the edges.
@@
Our baseline algorithms are SAT (Goh et al.,, 2022), SCNN (Yang et al., 2022a, ), and Simplicial MPNN (Bodnar et al.,, 2021). We will follow the batch training procedure seen in SAT (Goh et al.,, 2022): 40 trajectories are batched together to formulate a multidimensional simplicial data/ Each experiment will consist of 5000 iterations for all tested algorithms. For SAT, SCNN, and Bi-SCNN, we will be running two different architectural settings: 2 simplicial layers and 3 simplicial layers. As suggested by (Goh et al.,, 2022), all the simplicial layers are followed by a mean pooling layer, then 2 MLP readout layers with 30 channels, and finally a softmax layer. For each architectural setting, there will be three different activation functions: identity (id) function, leaky ReLu (LR) function, and tanh function. The SAT suggests the choices of Id and Tanh functions because odd activation functions lead to orientation-equivariant output, which is shown to perform well on a synthetic ocean drifter dataset (Goh et al.,, 2022). The classification accuracy is shown in Table 4 and the time to complete the experiments is recorded in Table 5. The number of trainable parameters of each setting is shown in Table 6. All the tested algorithms will have 40 simplicial filters except the MPNN. Again, due to the fact that an MPNN with a similar hyperparameter setting to the SCNN surpassed the capability of our hardware. For the classification accuracy of the MPNN, we can refer to the original paper (Bodnar et al.,, 2021) as it implemented a 4-layer MPNN with batch size 64 and filter size 64 on this exact dataset for the three tested activation functions in this experiment. We will a 2-layer MPNN with 1 filter only to measure the runtime of the experiment.
@@
In the Ocean Drifter Dataset, we can observe again that the Bi-SCNN has similar classification accuracy as the SCNN and MPNN for all three choices of action functions. Considering the fact that Bi-SCNN has the smallest parameter size among Bi-SCNN, SCNN, and MPNN, it is safe to say that the Bi-SCNN is more efficient. As for SAT, the reason Bi-SCNN is capable of consistently outperforming SAT is because the Bi-SCNN utilizes the Hodge decomposition: Bi-SCNN optimizes parameters for upper and lower adjacencies separately but the SAT uses the standard Hodge Laplacian. Additionally, the processing of the harmonic component of the input simplicial features is absent in SAT, making its simplicial embedding less representative than convolution-based algorithms.
One may notice that the run time reduction of Bi-SCNN is not as prominent in the ocean drifter complex as in the citation complex. This is because when we recorded the run time, the global pooling and the 2 MLP layers also contributed to the time it took to run the simplicial layers. The Bi-SCNN by itself does not optimize the global pooling and the 2 MLP layers, so these layers will perform indifferently among Bi-SCNN, SAT, and SCNN. Still, Bi-SCNN is the fastest among Bi-SCNN, SAT, SCNN, and MPNN when all three are compared under the same number of simplicial layers.
Model Trainable Parameters
SAT-2 5762
SAT-3 6722
SCNN-2 4712
SCNN-3 10142
Bi-SCNN-2 3750
Bi-SCNN-3 6450
MPNN-2 (ours) 167337644
MPNN (Bodnar et al.,, 2021) OOM
Table 6: The number of trainable parameters for each model on the Ocean Drifter dataset.
@@ 6 Conclusion
The Bi-SCNN was developed from a combination of simplicial convolution with weighted binary-sign propagation, effectively incorporating the upper and lower adjacencies to capture high-order structured simplicial features. The nonlinearity of Bi-SCNN is naturally introduced by feature normalization and binarization, allowing time-efficient operation of the Bi-SCNN. Bi-SCNN outperforms previously proposed SNN algorithms on citation and ocean drifter complexes in terms of time efficiency, while maintaining comparable prediction accuracy.